40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 23 13:30:06 2012, Jenne, Update, Green Locking, Yarm ALS laser is funny / dying 
    Reply  Mon Jul 23 18:01:30 2012, Jenne, Update, Green Locking, Yarm ALS laser is funny / dying 
    Reply  Thu Mar 14 16:41:48 2013, Jenne, Update, Green Locking, Yarm ALS laser is funny / dying 
       Reply  Thu Mar 14 17:31:57 2013, Manasa, Update, Green Locking, X arm green locked in TEM00 
          Reply  Fri Mar 15 01:57:02 2013, Jenne, Update, Green Locking, X arm green locked in TEM00 
       Reply  Fri Mar 15 02:19:01 2013, Jenne, Update, Green Locking, Yarm ALS laser is funny / dying 
          Reply  Fri Mar 15 15:26:13 2013, Koji, Update, Green Locking, Yarm ALS laser is funny / dying 
             Reply  Mon Mar 18 14:58:15 2013, Steve, Update, Green Locking, repair or replace dead NPRO 
                Reply  Thu Mar 28 07:58:15 2013, Steve, Update, Green Locking, NPRO repair 
Message ID: 8301     Entry time: Fri Mar 15 15:26:13 2013     In reply to: 8300     Reply to this: 8307
Author: Koji 
Type: Update 
Category: Green Locking 
Subject: Yarm ALS laser is funny / dying 

I took a look at the laser. It is probably the LD TEC (DTEC) failure.
As the temperature of the LD (DTMP) gradually deviated from 25degCish,
the DTEC voltage also went up from 2Vish to 2.1, 2.2... 

When DTEC reaches 3V, it stopped lasing. This cools the diode a bit, and
it start lasing but repeat the above process.

I am not sure which of the head and controller has the issue.

The situation did not improve much by reducing the pumping current (ADJ: -15).

BTW, Turning on/off the noise eater did not change the situation.

I think the head/controller set should be sent out to JDSU and find how they will say.

ELOG V3.1.3-